# ACE : Off-Policy Actor-Critic with Causality-Aware Entropy Regularization

### v1

@@ 1 Introduction
Figure 2: Motivating example. This task involves a robotic arm hammering a screw into a wall. normal-$\bullet$ Initially, the robotic arm approaches the desk moving on the z-axis and struggles with torque grasping, making z-axis positioning normal-$\uparrow$ and torque exploration normal-$\uparrow$ a priority. normal-$\blacktriangle$ As the training advances, the agent’s focus shifts to optimizing movement, prioritizing end-effector position (x-axis normal-$\uparrow$ and y-axis normal-$\uparrow$). normal-$\bigstar$ Finally, potential improvements lie in the stable and swift hammering, shifting focus back to torque normal-$\uparrow$ and placing down the object normal-$\uparrow$. The evolving causal weights, depicted on the left, reflect these changing priorities. See more examples in Appendix D.
@@ 2 Preliminary
Reinforcement Learning (RL). Within the standard framework of the Markov decision process (MDP), RL can be formulated as $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle$. Here, ${\cal S}$ denotes the state space, ${\cal A}$ the action space, $r:{\cal S\times A}\in[-R_{max},R_{max}]$ the reward function, and $\gamma\in(0,1)$ the discount factor, and $P(\cdot\mid s,a)$ stands for transition dynamics. The objective of an RL agent is to learn an optimal policy $\pi$ that maximizes the expected discounted sum of rewards, formulated as $\mathbb{E}\pi[\sum{t=0}^{\infty}\gamma^{t}r_{t}]$. For any $s\in\mathcal{S}$ and action $a\in\mathcal{A}$, the value of action $a$ under state $s$ is given by the action-value function $Q^{\pi}(s,a)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t})]$. In deep RL, the policy $\pi$ and the value function $Q$ are represented by neural network function approximations.
@@ 3.1 Causal Discovery on Policy-Reward Relationship
$r_{t}=r_{\mathcal{M}}\left({\mathbf{B}}_{{{\mathbf{s}}\to r|{\mathbf{a}}}}\odot{\mathbf{s}}_{t},{\color[rgb]{0.84,0.17,0.11}\definecolor[named]{pgfstrokecolor}{rgb}{0.84,0.17,0.11}\pgfsys@color@rgb@stroke{0.84}{0.17}{0.11}\pgfsys@color@rgb@fill{0.84}{0.17}{0.11}{\mathbf{B}}_{{\mathbf{a}}\to r|{\mathbf{s}}}}\odot{\mathbf{a}}_{t},\epsilon_{t}\right),$ (1)
@@
Suppose ${\mathbf{s}}_{t}$, ${\mathbf{a}}_{t}$, and $r_{t}$ follow the MDP model with Eq.(1). Under the Markov condition and faithfulness assumption, the structural vectors ${\mathbf{B}}_{{\mathbf{a}}\to r|{\mathbf{s}}}$ are identifiable.
@@ 3.2 Causality-aware Bellman Operator
$+\alpha{\color[rgb]{0.84,0.17,0.11}\definecolor[named]{pgfstrokecolor}{rgb}{0.84,0.17,0.11}\pgfsys@color@rgb@stroke{0.84}{0.17}{0.11}\pgfsys@color@rgb@fill{0.84}{0.17}{0.11}\mathcal{H}_{c}(\pi({\mathbf{a}}_{t+1}|{\mathbf{s}}_{t+1}))}]].$
@@ 3.3 Gradient-dormancy-guided Reset
In Figure 3, we depict the occurrence of the gradient dormancy phenomenon at the initial learning stages of SAC. However, in algorithms with superior sample efficiency like CausalSAC and ACE , there is a notable reduction in the dormancy degree during training. Particularly, when addressing challenging tasks such as pick place walls, which were previously unsolved by existing baselines, dormancy degrees persist at high levels without reset interventions in SAC and CausalSAC. We consider that dormancy degrees may impact sample efficiency and employ a soft reset method (Xu et al., 2023; Ash & Adams, 2020) to further decrease dormancy degree by periodically perturbing the policy network and critic network with a reset factor $\eta$, representing the magnitude of weight resetting:
@@
Combining the causality-aware entropy regularization and gradient-dormancy-guided reset mechanism, we propose our algorithm ACE : off-policy Actor-critic with Causality-aware Entropy regularization. The pseudocode and the implementation details are provided in Appendix E.
@@ 4.1 Evaluation on various benchmark suites
Figure 5: Locomotion tasks. Average return of ACE  SAC, TD3 on locomotion tasks provided by MuJoCo and DMControl benchmark suites. Solid curves depict the mean of six trials, and shaded regions correspond to the one standard deviation. See Figure 20 and 21 in the Appendix for an overall comparison of locomotion tasks.
@@ 4.2 Ablation Studies
We observe that all components of our proposed methodology significantly contribute to the final performance of ACE . Notably, our causal-weighted entropy introduced in CausalSAC exhibits the most pronounced enhancement compared to SAC, emphasizing the effectiveness of causality-aware exploration. Simultaneously, our reset mechanism also plays a crucial role in further improving performance, especially in sparse reward tasks. Interestingly, applying the gradient-dormancy-guided reset solely to SAC demonstrates improved performance and sample efficiency, suggesting that the gradient dormant phenomenon might indeed be one of the factors contributing to SAC’s suboptimal performance in hard tasks.
@@ 6 Conclusions and Discussion
Looking ahead, there are several directions for further exploration stemming from this work. First, despite our analysis of computing causal weights in the latent space and their application to DrM (Xu et al., 2023) in Appendix J, there is untapped potential for exploring additional applications of causality-aware exploration to enhance the sample efficiency of visual RL. Second, it is worth discussing the prospect of leveraging dynamic models in model-based RL to facilitate more efficient computation of causality between rewards and policy. Concerning time efficiency, a detailed analysis is provided in Appendix K, confirming the high efficiency and cost-effectiveness of our algorithm. We believe that the implications of causality-aware exploration extend beyond our current work, offering more effective solutions to enhance the sample efficiency of online RL.
@@ Appendix B Extensive Related Works
Inspired by human cognition, dividing the original task into multiple stages for policy learning has been explored in RL through heuristic methods tailored to specific tasks (Jinnai et al., 2019; Liu et al., 2023). For instance, hierarchical RL (Sutton et al., 1999; Pateria et al., 2021) has implemented this idea by introducing additional subgoal spaces or leveraging the semi-Markov assumption. This allows the agent to segment tasks into different subtasks (Nachum et al., 2018), options (Machado et al., 2023), or skills (Guan et al., 2022) as multiple stages, enabling policies to possess distinct exploration capabilities across various stages, ultimately enhancing the success of agents in complex continuous control tasks. In contrast to explicit multi-stage approaches, our proposed method ACE  does not necessitate a clear task stage division. Instead, it can identify and prioritize actions with a high potential for reward through causal discovery dynamically, emphasizing the various importance of each action dimension in different stages for enhanced exploration and performance.
@@ E.1 Algorithm instantiation
$\mathcal{T}_{c}^{\pi}Q({\mathbf{s}}_{t},{\mathbf{a}}_{t})\triangleq r({\mathbf{s}}_{t},{\mathbf{a}}_{t})+\gamma\mathbb{E}_{{\mathbf{s}}_{t+1}\sim P}\left[\mathbb{E}_{{\mathbf{a}}_{t}\sim\pi}[Q({\mathbf{s}}_{t+1},{\mathbf{a}}_{t+1})+\alpha{\color[rgb]{0.84,0.17,0.11}\definecolor[named]{pgfstrokecolor}{rgb}{0.84,0.17,0.11}\pgfsys@color@rgb@stroke{0.84}{0.17}{0.11}\pgfsys@color@rgb@fill{0.84}{0.17}{0.11}\mathcal{H}_{c}(\pi({\mathbf{a}}_{t+1}|{\mathbf{s}}_{t+1}))}]\right].$ (7)
@@
     // Causal discovery   if every $I$ environment step then      Sample all $N_{c}$ transitions from local buffer $\mathcal{D}_{c}$      Update causal weight matrix ${\mathbf{B}}_{{\mathbf{a}}\rightarrow r|{\mathbf{s}}}$   end if
     // Policy optimization   for each gradient step do      Sample $N$ transitions $({\mathbf{s}},{\mathbf{a}},r,{\mathbf{s}}^{\prime})$ from $\mathcal{D}$      Compute causality-aware entropy $\mathcal{H}_{c}(\pi(\cdot|{\mathbf{s}}))$      Calculate the target $Q$ value      Update $Q_{\phi}$ by $\min_{\phi}\left(\mathcal{T}_{c}Q_{\phi}-Q_{\phi}\right)^{2}$      Update $\pi_{\theta}$ by $\max_{\theta}Q_{\phi}(s,a)$   end for
     // Reset mechanism   if every reset interval then      Calculate the gradient dormant degree $\beta_{\gamma}$      Initialize a random network with weights $\phi_{i}$      Soft reset $\pi_{\theta}$ by $\theta_{t}=(1-\eta)\theta_{t-1}+\eta\phi_{i}$      Soft reset $Q_{\phi}$ by $\phi_{t}=(1-\eta)\phi_{t-1}+\eta\phi_{i}$      Reset the state of the policy optimizer and Q optimizer   end if
@@ E.2 Hyperparameters
$Q$-value network
MLP with hidden size 512
$V$-value network
MLP with hidden size 512
policy network
Gaussian MLP with the hidden size 512
discounted factor $\gamma$
0.99
soft update factor $\tau$
0.005
learning rate $\alpha$
0.0003
batch size $N$
512
policy updates per step
1
value target updates interval
2
sample size for causality $N_{c}$
10000
causality computation interval $I$
10000
max reset factor $\alpha_{max}$
0.8
reset interval
200000
dormancy threshold $\tau$
0.025
@@ Appendix I Generalizability of ACE
The proposed causality-aware entropy and gradient-dormancy-guided reset mechanism are versatile and effective plug-ins, thus they could be intergrated into various backbone Max-Entropy algorithms.
We integrate ACE into a recent BAC algorithm, which is effective in challenging failure-prone tasks. Results on the challenging dexterous hand manipulation tasks show that the ad-hoc ACE-BAC algorithm could outperform the BAC algorithm, refer to Figure 25. These observations demonstrate the effectiveness and generalizability of our proposed mechanisms and also shed light on the further applications of our mechanisms in enhancing existing RL algorithms.
@@
Appendix K Computing Infrastructure and Computational Time

### v3

@@ 1 Introduction
Figure 2: Motivating example. This task involves a robotic arm hammering a screw into a wall. $\bullet$ Initially, the robotic arm approaches the desk moving on the z-axis and struggles with torque grasping, making z-axis positioning $\uparrow$ and torque exploration $\uparrow$ a priority. $\blacktriangle$ As the training advances, the agent’s focus shifts to optimizing movement, prioritizing end-effector position (x-axis $\uparrow$ and y-axis $\uparrow$). $\bigstar$ Finally, potential improvements lie in the stable and swift hammering, shifting focus back to torque $\uparrow$ and placing down the object $\uparrow$. The evolving causal weights, depicted on the left, reflect these changing priorities. See more examples in Appendix D.
@@ 2 Preliminary
Reinforcement Learning (RL). Within the standard framework of the Markov decision process (MDP), RL can be formulated as $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle$. Here, ${\cal S}$ denotes the state space, ${\cal A}$ the action space, $r:{\cal S\times A}\in[-R_{max},R_{max}]$ the reward function, and $\gamma\in(0,1)$ the discount factor, and $P(\cdot\mid s,a)$ stands for transition dynamics. The objective of an RL agent is to learn an optimal policy $\pi$ that maximizes the expected discounted sum of rewards, formulated as $\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$. For any $s\in\mathcal{S}$ and action $a\in\mathcal{A}$, the value of action $a$ under state $s$ is given by the action-value function $Q^{\pi}(s,a)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t})]$. In deep RL, the policy $\pi$ and the value function $Q$ are represented by neural network function approximations.
@@ 3.1 Causal Discovery on Policy-Reward Relationship
$r_{t}=r_{\mathcal{M}}\left({\mathbf{B}}_{{{\mathbf{s}}\to r|{\mathbf{a}}}}\odot{\mathbf{s}}_{t},{\color[rgb]{0.84,0.17,0.11}\definecolor[named]{pgfstrokecolor}{rgb}{0.84,0.17,0.11}{\mathbf{B}}_{{\mathbf{a}}\to r|{\mathbf{s}}}}\odot{\mathbf{a}}_{t},\epsilon_{t}\right),$ (1)
@@
Suppose $s_{t}$, $a_{t}$, and $r_{t}$ follow the MDP model with no unobserved confounders, as in Eq.(1). Under the Markov condition and faithfulness assumption, the structural vectors ${\mathbf{B}}_{{\mathbf{a}}\to r|{\mathbf{s}}}$ are identifiable.
Note that such a theorem guarantees the identifiability of the correct graph in an asymptotic manner. Additionally, by imposing further assumptions on the data generation mechanism, we could uniquely identify the causal effects. See the proof in Appendix A.1 for details.
@@ 3.2 Causality-aware Bellman Operator
$+\alpha{\color[rgb]{0.84,0.17,0.11}\definecolor[named]{pgfstrokecolor}{rgb}{0.84,0.17,0.11}\mathcal{H}_{c}(\pi({\mathbf{a}}_{t+1}|{\mathbf{s}}_{t+1}))}]].$
@@ 3.3 Gradient-dormancy-guided Reset
(a) In simpler tasks, CausalSAC performs adequately, with observed gradient dormancy degrees decreasing to near-zero. This indicates that in less complex environments, the issue of gradient dormancy becomes negligible.
(b) In challenging tasks, both SAC and CausalSAC exhibit high dormancy degrees without achieving expected performance. The gradient dormancy-guided reset mechanism in ACE, however, effectively lowers the gradient dormancy degree.
@@
In Figure 3, we depict the occurrence of the gradient dormancy phenomenon at the initial learning stages of SAC. However, in algorithms with superior sample efficiency like CausalSAC and ACE , there is a notable reduction in the dormancy degree during training. In tasks that CausalSAC can effectively solve, the gradient dormancy degree can decrease to near-zero levels, as shown in Figure 3. Particularly, when addressing challenging tasks such as pick place walls, which were previously unsolved by existing baselines, dormancy degrees persist at high levels without reset interventions in SAC and CausalSAC. Hence, we speculate that this may represent a potential local optimum for causal-aware exploration. We consider that dormancy degrees may impact sample efficiency and employ a soft reset method (Xu et al., 2023; Ash & Adams, 2020) to further decrease dormancy degree by periodically perturbing the policy network and critic network with a reset factor $\eta$, representing the magnitude of weight resetting:
@@
3.4 Algorithm instantiation
Combining the causality-aware entropy regularization and gradient-dormancy-guided reset mechanism, we propose our algorithm ACE : off-policy Actor-critic with Causality-aware Entropy regularization. The pseudocode and further implementation details are provided in Appendix E.
To effectively compute $\mathbb{B}_{a_{i}\rightarrow r|s}$, we adopt the well-regarded DirectLiNGAM (Shimizu et al., 2011) method. The main implementation idea of training DirectLiNGAM is as follows. In the first phase, it estimates a causal ordering for all variables of interest (i.e., state, action, and reward variables) based on the independence and non-Gaussianity characteristics of the root variable. The causal ordering is a sequence that implies the latter variable cannot cause the former one. In the second phase, DirectLiNGAM estimates the causal effects between variables using some conventional covariance-based methods. Besides, we formulate a training regime wherein we iteratively adjust the causal weights for the policy at regular intervals $I$ on a local buffer $\mathcal{D}_{c}$ with fresh transitions to reduce computation cost.
Given the causal weight matrix ${\mathbf{B}}_{{\mathbf{a}}\rightarrow{\mathbf{r}}|{\mathbf{s}}}$, we could obtain the causality-aware entropy $\mathcal{H}_{c}(\pi(\cdot|{\mathbf{s}}))$ through Eq.(2). Based on the causality-aware entropy, then the $Q$-value for a fixed policy $\pi$ could be computed iteratively by applying $\mathcal{T}_{c}^{\pi}$. Based on the policy evaluation, we can adopt many off-the-shelf policy optimization oracles; we chose SAC as the backbone technique primarily for its simplicity in our primary implementation of CausalSAC and ACE.
For each reset interval, we calculate the gradient dormancy degree, initialize a random network with weights $\phi_{i}$, and soft reset the policy network $\pi_{\theta}$ and the Q network $Q_{\phi}$.
@@ 4.1 Evaluation on various benchmark suites
Figure 5: Locomotion tasks. Average return of ACE  SAC, TD3 on locomotion tasks provided by MuJoCo and DMControl benchmark suites. Solid curves depict the mean of six trials, and shaded regions correspond to the one standard deviation. See Figure 23 and 24 in the Appendix for an overall comparison of locomotion tasks.
@@ 4.2 Ablation Studies
Figure 9: Ablation studies on SOTA exploration techniques. Learning curves of ACE, CausalSAC, SAC, and SOTA exploration technique RND on three sparse reward tasks.
Refer to captionRefer to caption
Figure 11: Ablation on reset mechanism factors. Learning curves of ACE w/wo fixed reset factor $\eta$ and fixed initializer $\Phi$.
@@
We observe that all components of our proposed methodology significantly contribute to the final performance of ACE . Even without the reset mechanism, CausalSAC (vanilla ACE) demonstrates superior performance. Further, results on sparse reward tasks show that CausalSAC outperforms other SOTA exploration techniques. This highlights its potential in challenging environments.
Simultaneously, our reset mechanism also plays a crucial role in further improving performance, especially in sparse reward tasks. Interestingly, applying the gradient-dormancy-guided reset solely to SAC demonstrates improved performance and sample efficiency, suggesting that the gradient dormant phenomenon might indeed be one of the factors contributing to SAC’s suboptimal performance in hard tasks.
@@
Ablation studies on reset mechanism factors. We conduct ablation studies on the fixed reset factor $\eta$ and the fixed network initializer $\phi$ across three tasks. As shown in Figure 11, fixing the reset factor (i.e., no longer guiding it by the dormancy degree) results in a significant performance drop. This highlights the importance of the proposed dormancy degree in our method. On the other hand, fixing the network initializer $\phi$ does not affect the performance of the dormancy-guided reset mechanism in ACE.
@@ 6 Conclusions and Discussion
Looking ahead, there are several directions for further exploration stemming from this work. First, despite our analysis of computing causal weights in the latent space and their application to DrM (Xu et al., 2023) in Appendix J, there is untapped potential for exploring additional applications of causality-aware exploration to enhance the sample efficiency of visual RL. Second, it is worth discussing the prospect of leveraging dynamic models in model-based RL to facilitate more efficient computation of causality between rewards and policy. Concerning time efficiency, a detailed analysis is provided in Appendix L, confirming the high efficiency and cost-effectiveness of our algorithm. We believe that the implications of causality-aware exploration extend beyond our current work, offering more effective solutions to enhance the sample efficiency of online RL.
@@ Appendix B Extensive Related Works
Inspired by human cognition, dividing the original task into multiple stages for policy learning has been explored in RL through heuristic methods tailored to specific tasks (Jinnai et al., 2019; Liu et al., 2023). For instance, hierarchical RL (Sutton et al., 1999; Pateria et al., 2021) has implemented this idea by introducing additional subgoal spaces or leveraging the semi-Markov assumption. This allows the agent to segment tasks into different subtasks (Nachum et al., 2018), options (Machado et al., 2023), or skills (Guan et al., 2022) as multiple stages, enabling policies to possess distinct exploration capabilities across various stages, ultimately enhancing the success of agents in complex continuous control tasks. Further, an interesting work is Human-AI shared control via Policy Dissection (Li et al., 2022). While both ACE and Policy Dissection draw inspiration from neuroscience, particularly in the realm of motion primitives, they operate in different paradigms. Policy Dissection fosters human-AI shared control, enabling human collaboration with RL agents in intricate environments. It establishes an interpretable interface on the agent, allowing humans to directly influence agent behavior and facilitate shared control.
In contrast to explicit multi-stage approaches or human-guided task stage division, our proposed method ACE  does not necessitate a clear task stage division. Instead, it can identify and prioritize actions with a high potential for reward through causal discovery dynamically, emphasizing the various importance of each action dimension in different stages for enhanced exploration and performance.
@@ Appendix D More Examples of Causal Weights
Figure 21: Visualizations of causal weights in Adroit Door and Pen tasks (29 action dimensions). These curves demonstrate how causal weights control individual finger joints alongside the entire arm, with detailed explanations provided. Notably, the crucial finger joints differ across tasks. This clearly highlights the effectiveness of our causal graph calculation in high-dimensional settings involving multi-object control.
@@ E.1 Algorithm instantiation
$\mathcal{T}_{c}^{\pi}Q({\mathbf{s}}_{t},{\mathbf{a}}_{t})\triangleq r({\mathbf{s}}_{t},{\mathbf{a}}_{t})+\gamma\mathbb{E}_{{\mathbf{s}}_{t+1}\sim P}\left[\mathbb{E}_{{\mathbf{a}}_{t}\sim\pi}[Q({\mathbf{s}}_{t+1},{\mathbf{a}}_{t+1})+\alpha{\color[rgb]{0.84,0.17,0.11}\definecolor[named]{pgfstrokecolor}{rgb}{0.84,0.17,0.11}\mathcal{H}_{c}(\pi({\mathbf{a}}_{t+1}|{\mathbf{s}}_{t+1}))}]\right].$ (7)
@@
   {mdframed} [linewidth=0.8pt, linecolor=myblue]   // Causal discovery   if every $I$ environment step then      Sample all $N_{c}$ transitions from local buffer $\mathcal{D}_{c}$      Update causal weight matrix ${\mathbf{B}}_{{\mathbf{a}}\rightarrow r|{\mathbf{s}}}$   end if
   {mdframed} [linewidth=0.8pt, linecolor=mygreen]   // Policy optimization   for each gradient step do      Sample $N$ transitions $({\mathbf{s}},{\mathbf{a}},r,{\mathbf{s}}^{\prime})$ from $\mathcal{D}$      Compute causality-aware entropy $\mathcal{H}_{c}(\pi(\cdot|{\mathbf{s}}))$      Calculate the target $Q$ value      Update $Q_{\phi}$ by $\min_{\phi}\left(\mathcal{T}_{c}Q_{\phi}-Q_{\phi}\right)^{2}$      Update $\pi_{\theta}$ by $\max_{\theta}Q_{\phi}(s,a)$   end for
   {mdframed} [linewidth=0.8pt, linecolor=myyellow]   // Reset mechanism   if every reset interval then      Calculate the gradient dormant degree $\beta_{\gamma}$      Initialize a random network with weights $\phi_{i}$      Soft reset $\pi_{\theta}$ by $\theta_{t}=(1-\eta)\theta_{t-1}+\eta\phi_{i}$      Soft reset $Q_{\phi}$ by $\phi_{t}=(1-\eta)\phi_{t-1}+\eta\phi_{i}$      Reset the state of the policy optimizer and Q optimizer   end if
@@ E.2 Hyperparameters
$Q$-value network MLP with hidden size 512
$V$-value network MLP with hidden size 512
policy network Gaussian MLP with the hidden size 512
discounted factor $\gamma$ 0.99
soft update factor $\tau$ 0.005
learning rate $\alpha$ 0.0003
batch size $N$ 512
policy updates per step 1
value target updates interval 2
sample size for causality $N_{c}$ 10000
causality computation interval $I$ 10000
max reset factor $\alpha_{max}$ 0.8
reset interval 200000
dormancy threshold $\tau$ 0.025
@@ Appendix I Generalizability of ACE
The proposed causality-aware entropy and gradient-dormancy-guided reset mechanism are versatile and effective plug-ins, thus they could be integrated into various RL backbone algorithms and techniques.
@@
Figure 29: Performance of ACE with higher-UTD methods. We applied ACE to REDQ by integrating causality-aware entropy and a dormancy-guided reset mechanism. The results show that our method significantly improves REDQ’s sample efficiency and performance in multiple tasks. Solid curves depict the mean of 6 trials, and shaded regions correspond to the one standard deviation.
halfcheetah-medium-v2 hopper-medium-v2 walker2d-medium-v2 maze2d-medium-v1
CQL 47.04 $\pm$ 0.22 59.08 $\pm$ 3.77 80.75 $\pm$ 3.28 86.11 $\pm$ 9.68
CQL+ACE 50.23 $\pm$ 0.43 60.62 $\pm$ 3.97 85.27 $\pm$ 3.68 89.61 $\pm$ 18.72
Table 2: Performance comparison of CQL and CQL+ACE across various tasks.
Compatibility with Max-Entropy algorithms.
We integrate ACE into a recent BAC algorithm, which is effective in challenging failure-prone tasks. Results on the challenging dexterous hand manipulation tasks show that the ad-hoc ACE-BAC algorithm could outperform the BAC algorithm, refer to Figure 28.
Compatibility with high-UTD techniques.
Higher-UTD (Update-to-Data) methods are popular and effective in improving the sample efficiency and performance of reinforcement learning algorithms. We integrate ACE into the popular REDQ (Chen et al., 2020) algorithm, which is known for its efficiency in continuous control tasks. Results on the Walker2d and CheetahRun benchmark tasks, shown in Figure 29 demonstrate that ACE significantly enhances REDQ’s sample efficiency and overall performance across a range of tasks.
Compatibility with offline RL setting.
Excitingly, we observed that CQL with an ACE backbone outperforms CQL. This not only underscores the effectiveness and versatility of ACE but also highlights a promising direction for future exploration.
These observations demonstrate the effectiveness and generalizability of our proposed mechanisms and also shed light on the further applications of our mechanisms in enhancing existing RL algorithms.
@@
Appendix K The effectiveness of ACE in solving hard exploration problems.
The effectiveness of ACE is particularly evident in scenarios where the causal relationship between action and reward, conditioned on the state, can be misleading and hurt performance.
Addressing Dummy Actions with the Reset Mechanism.
For scenarios involving dummy actions, the resetting mechanism helps to solve it. To be specific, such dummy actions would trap the exploration in local optima and hence the gradient dormant would be high (i.e., low policy gradients, high gradient dormancy degree). Then our reset mechanism intervenes by perturbing the network, encouraging unbiased exploration.
A concrete example is the pick-and-place-wall task, where the wall initially biases the agents’ exploration towards vertical and grasping actions. In contrast to the high dormancy degree and poor performance observed with all other baselines, ACE overcomes this challenge with its dormancy-guided reset mechanism, as evidenced by the dormancy degree curves in Figure 3 (y-axis=dormancy degree) and the performance data in Figure 4.
Handling Misleading or Imperfect Rewards.
For scenarios with misleading or imperfect rewards, we could simply turn it into a sparse reward setting. Our experiments, as shown in Figures 7 and  27, reveal ACE’s effectiveness in sparse reward environments. This advantage primarily arises from ACE’s causal-aware entropy, which discerns the implicit action-reward relationship based on the state.
Furthermore, ACE’s versatility also allows it to be combined with alternative intrinsic reward approaches, enhancing its applicability across various scenarios. This adaptability ensures that specific challenges, such as those involving misleading rewards.
Appendix L Computing Infrastructure and Computational Time